If g(x) = 4|x-3|+2. Find g(1)

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39800)
Answer by ikleyn(53765)

.
Simply substitute x= 1 into the formula, and you will get

g(1) = 4|1-3|+2 = 4*|-2|+2 = 4*2 + 2 = 8 + 2 = 10.

Answer. g(1) = 10.